Took a Test the other day and got these 2 questions wrong i need help confused how to do them.
Simplify these 2 and show work so i can understand what to do with them later.
12^2/3 Possible answers are 2 3rd root of 6 4 4th root of 2
4 3rd root of 12 2 3rd root of 18

(4^2/3)^9/10 Possible answers are 5th root of 2 5th root of 4
2 5th root of 2 4 5th root of 4
hope this is enough info to help. Thank you for the help

Answer by edjones(7569)

12^(2/3)
=144^(1/3)
=(18*8)^(1/3)
=2(18)^(1/3)
.
Ed